if the acceptance cone for the given fiber is 68.16 calculate the maximum entrance angle and numerical aperture if the cladding glass has a refractive index of a 1.52 calculate the refractive index of cone


1
Expert's answer
2021-07-05T08:43:18-0400

"NA=n\sin \alpha=1.41,"

"\alpha_{max}=\alpha \sqrt{ NA}=80.94°,"

"n'=NA=1.41."